OpenStudy (anonymous):

find product and simplify 2x-8/ x-1 * 4x+8/ x-4

OpenStudy (anonymous):

\[\frac{ 2x-8 }{ x-1 }\times \frac{ 4x+8 }{ x-4 }\]

OpenStudy (jdoe0001):

\(\bf \cfrac{ 2x-8 }{ x-1 }\times \cfrac{ 4x+8 }{ x-4 }\implies \cfrac{ 2\cancel{ (x-4) } }{ x-1 }\times \cfrac{ 4x+8 }{ \cancel{ x-4 } }\) any ideas on what's left?